OpenStudy (anonymous):

18x over 45 how would u simplify that

OpenStudy (amistre64):

18x/45. factor 18 and 45 and cross out all the numbers that are the same. 18 = (2)(9) = (2)(3)(3) 45 = (3)(15) = (3)(3)(5) they have (3)(3)/(3)(3) in common so get rid of them because they simply equal the number 1. we are left with: (2)x/(5) = 2x/5